OpenStudy (anonymous):

I'm having trouble with equations that have more than one variable. I need to solve for "y" and I've gotten the problem from 6*y−4−3*y+1=12 to 6*y-3*y=15. First of all did I simplify right. Second, what is a formula for any equation with more than one variable.

OpenStudy (anonymous):

I forgot to say that (if it helps) I know "y"=5

OpenStudy (anonymous):

Yes, you simplified correctly. At this point, apply the distributive property in reverse:\[6y-3y=(6-3)y=3y\]

OpenStudy (anonymous):

I so sorry but you lost me.

OpenStudy (anonymous):

Would you agree with me that \((6-3)y = 6y-3y\) by the distributive property?

OpenStudy (anonymous):

Yes, that part makes sense.

OpenStudy (anonymous):

Oh now I understand. Thank you so much and have a happy national battery day

OpenStudy (anonymous):

So since we can agree \(6y-3y=(6-3)y=3y\), we can turn your equation \(6y-3y=12\) into \(3y=12\), which I am confident you can solve.

OpenStudy (anonymous):

Wait, I'm lost again how did you get 12.

OpenStudy (anonymous):

Sorry...typo. Meant 15 as you wrote above. We turn \(6y-3y=15\) into \(3y=15\).

OpenStudy (anonymous):

Ok give me a moment to analyze in my head.